OpenStudy (anonymous):

1. A company produces two varieties of pens: alpha and beta. Each alpha pen needs twice as much labour time as the beta pen. If only beta pens are manufactured, then the company can produce 500 pens per day. The market can take only up to 150 alpha pens and 250 beta pens per day from the company. The profits earned by the company on selling an alpha pen and a beta pen are Rs 8 and Rs 5 respectively. What are the respective numbers of alpha and beta pens that the company needs to produce in order to maximize the profit?

OpenStudy (compassionate):

I'll answer it if you give me a medal :)

OpenStudy (anonymous):

i dont know hw 2 giv medal and its procedure pls tell the sol

OpenStudy (compassionate):

I'm working on it.

OpenStudy (anonymous):

ok thank u

OpenStudy (compassionate):

See that blue thing that says 'best response.' click that. It's on the right of the screen.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pls try it pls tell the sol

OpenStudy (compassionate):

Don't think I forgot about you: 250 beta pens and 125 alpha pens. Then you get Rs. 50 profit.

OpenStudy (anonymous):

pls tell me procedure

OpenStudy (compassionate):

No. of max. beta pens that can be produced= 500 Therefore(sign), no. of max. alpha pens that can be made=no. of max. beta pens/2= 500/2= 250(given) To make max. profit, one type of pen should be sold at max. When selling max. beta pens, No. of beta pens sold = 250 No. of alpha pens sold=125(since no. of beta pens that can be made more = 250, therefore no of alpha pens made =250/2=125) Total profit=Rs.(250*5+125*8) = Rs.(1250+1000)=Rs. 2250 When selling max. alpha pens, No. of alpha pens sold=150 No. of beta pens sold= 200(since…..please write own) Total profit= Rs.(150*8 + 200*5)=Rs.(1200+1000)=Rs. 2200 Therefore(use sign), Profit when selling max. beta pens= Rs.(2250-2200)=Rs. 50 Therefore, the company has to sell 250 beta pens and 150 alpha pens in order to make max. profit

OpenStudy (anonymous):

If you want to know more about such problems. Google with keyword PRODUCT MIX problem. Basically it's an optimization problem.